The sum of two numbers is 39. Their difference is 11. What are the two numbers?
Answer by consc198(59)

Lets say that one of the numbers is a and the other one b
So a + b = 39 and a - b = 11
So a = 25 and b = 14
Lets substitute:
25 + 14 = 39
25 - 14 = 11

The two numbers are 25 and 11
For more maths help please send as an email to mathshelp@mathshomeworkhelp.4t.com or visit our website www.mathshomeworkhelp.4t.com